# Analysis of Linear Mode Connectivity via Permutation-Based Weight Matching: With Insights into Other Permutation Search Methods

### v4

@@ Abstract
Recently, Ainsworth et al. (2023) showed that using weight matching (WM) to minimize the $L_{2}$ distance in a permutation search of model parameters effectively identifies permutations that satisfy linear mode connectivity (LMC), where the loss along a linear path between two independently trained models with different seeds remains nearly constant. This paper analyzes LMC using WM, which is useful for understanding stochastic gradient descent’s effectiveness and its application in areas like model merging. We first empirically show that permutations found by WM do not significantly reduce the $L_{2}$ distance between two models, and the occurrence of LMC is not merely due to distance reduction by WM itself. We then demonstrate that permutations can change the directions of the singular vectors, but not the singular values, of the weight matrices in each layer. This finding shows that permutations found by WM primarily align the directions of singular vectors associated with large singular values across models. This alignment brings the singular vectors with large singular values, which determine the model’s functionality, closer between the original and merged models, allowing the merged model to retain functionality similar to the original models, thereby satisfying LMC. This paper also analyzes activation matching (AM) in terms of singular vectors and finds that the principle of AM is the same as that of WM. Finally, we analyze the difference between WM and the straight-through estimator (STE), a dataset-dependent permutation search method, and show that WM can be more advantageous than STE in achieving LMC among three or more models.
@@ 1 Introduction
Deep learning has significantly advanced various fields, including image classification, speech recognition, and natural language processing (Vaswani et al., 2017; van den Oord et al., 2016; Zhao et al., 2023). Large-scale neural networks (NNs) are widely used in these applications, and optimizing their parameters poses a massive non-convex optimization problem. Remarkably, stochastic gradient descent (SGD), which is widely used for training NNs, is known to find good solutions despite its simplicity. One hypothesis for this seemingly counterintuitive phenomenon is that the landscape of the loss function may be much simpler than previously thought. Several studies (Garipov et al., 2018; Draxler et al., 2018; Freeman & Bruna, 2017) have found that different neural network solutions can be connected by nonlinear paths with almost no increase in loss. Recently, Entezari et al. (2022) conjectured that 1.1 holds, considering all possible permutation symmetries of NNs:
@@
Here, the barrier represents the increase in loss observed when linearly interpolating between the weights of the two models. A sufficiently small barrier implies that the model obtained by linear interpolation of the two models performs the same as the original models. If the barrier between two models is sufficiently small, we say that linear mode connectivity (LMC) is satisfied between them (Frankle et al., 2020). 1.1 suggests that most SGD solutions can be transferred into the same loss basin using permutations. Indeed, some studies (Ainsworth et al., 2023; Singh & Jaggi, 2020) have experimentally demonstrated that this conjecture is valid for various datasets and models using weight matching (WM), which identifies permutations that minimize the $L_{2}$ distance between the weights of two models.
Understanding LMC principles based on permutation symmetries is important not only for comprehending how SGD works in deep learning but also for its application in model merging (Singh & Jaggi, 2020), where two independently trained models are combined. The method of finding permutations using only the $L_{2}$ distance is particularly versatile, dataset-independent, and computationally efficient. In fact, several studies (Singh & Jaggi, 2020; Wang et al., 2020; Peña et al., 2023) have proposed applications of permutation symmetries in model merging, federated learning, and continual learning.
The current theoretical analysis of LMC depends on the feasibility of closely matching NN weights using permutations. Recently, Zhou et al. (2023) proved that if the distance between the weights of two models can be sufficiently reduced through permutation, then LMC holds. Intuitively, given two SGD solutions $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$, if a permutation $\pi$ can be found such that $\bm{\theta}_{a}\approx\pi(\bm{\theta}_{b})$, it is reasonable to expect that the outputs of the interpolated model will approximately equal the outputs of the original models $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$.
However, this paper shows that this intuitive explanation is inaccurate and clarifies a more fundamental reason why LMC holds through permutations found by WM. Our analysis shows that even if LMC holds, the permutations found by WM do not significantly reduce the distance between the two models (at most about a 20% reduction). We then explain this in terms of singular vectors with large singular values. Our analysis not only reveals the principle behind WM but also shows that WM may be more advantageous in merging more than three models compared to other methods such as STE.
@@
1. Demonstrating that the $L_{2}$ distance reduced by WM is not the direct cause of LMC.
This paper empirically shows that permutations found by WM do not significantly reduce the $L_{2}$ distance between the two models. Our results indicate that, even when LMC is satisfied, permutations reduce the model weight distance by no more than 20%. Supported by a Taylor approximation, our findings suggest that reducing the $L_{2}$ distance through permutations is not the direct reason for LMC satisfaction.
@@
We provide evidence that WM satisfies LMC by aligning the directions of singular vectors with large singular values in each layer’s weights. This alignment ensures that the singular vectors with large singular values, which determine the model’s functionality, become similar between the merged and original models. Additionally, we show that, from the perspective of the input distribution at each hidden layer, aligning singular vectors with large singular values can efficiently approximate the functionality between two models, even if the $L_{2}$ distance cannot be significantly reduced. As a result, the merged model retains functionality similar to the original models, which facilitates LMC. We also conducted experiments with AM and found that the reason why LMC holds in AM is the same as in WM.
@@
To distinguish WM from other permutation search methods that are independent of $L_{2}$ distance, we examine the straight-through estimator (STE), which focuses on minimizing the barrier itself rather than the $L_{2}$ distance. Our experiments reveal that the permutations found by STE do not align the directions of singular vectors, which is a critical difference compared to WM in achieving LMC. Furthermore, we demonstrate experimentally that this difference significantly impacts the satisfaction of LMC among three or more models.
@@ 2.1 Notation
For any natural number $k\in\mathbb{N}$, let $[k]=\{1,2,\dots,k\}$. Bold uppercase variables represent tensors, including matrices (e.g., $\bm{X}$), and bold lowercase variables (e.g., $\bm{x}$) represent vectors. For any tensor $\bm{X}$, its vectorization is denoted by $\textrm{vec}(\bm{X})$, and $\|\bm{X}\|$ denotes its Frobenius ($L_{2}$) norm.
@@ 2.3 Linear Mode Connectivity (LMC)
Intuitively, the barrier represents the increase in loss because of the linear interpolation of the two models. Two models $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$ are said to be linearly mode connected when the loss barrier between them is approximately zero.
@@ 2.4 Permutation Selection
In WM, we search for a permutation that minimizes the $L_{2}$ distance between two models [^1] Although only the weights are considered here, the biases can also be dealt with by concatenating the biases and the weights.:
@@
where, without loss of generality, let $\bm{P}_{L}=\bm{I}$ and $\bm{P}_{0}=\bm{I}$, and $\bm{I}$ is an identity matrix. This minimization problem is known as the sum of the bilinear assignments problem, which is NP-hard (Koopmans & Beckmann, 1957; Sahni & Gonzalez, 1976; Ainsworth et al., 2023). Recently, Peña et al. (2023) proposed solving Equation 2 using Sinkhorn’s algorithm (Adams & Zemel, 2011) by considering it as an optimal transport problem. We adopt their method because it allows for the optimization of all layers simultaneously, unlike the method by Ainsworth et al. (2023), and potentially finds better solutions.
@@ 3 Motivating Observations
Previous studies have suggested that the closeness of two parameters in terms of $L_{2}$ distance is important for satisfying LMC. For example, Zhou et al. (2023) showed that LMC holds if a commutativity property is satisfied. This property holds if, for all layers $\ell$, $\bm{W}^{(a)}_{\ell}-\bm{P}_{\ell}\bm{W}^{(b)}_{\ell}\bm{P}_{\ell-1}^{\top}=\bm{0}$. Zhou et al. (2023) argued in Section 5.2 that since WM finds the permutation that minimizes Equation 2, WM can be seen as searching for permutations that satisfy the commutativity property. In particular, there is a huge number of permutations because the total number of possible permutations grows exponentially as the number of layers and the width increase, and thus, some of them may sufficiently reduce the distance between the two models. However, this section explains from the perspective of a Taylor approximation that this intuition is not always correct. Our results demonstrate that even when LMC is satisfied, the permutations found by WM do not necessarily bring the models as close as expected. The facts observed from the experiments in this section motivate us to explore other reasons for satisfying LMC in the following sections.
@@ 3.1 Closeness of Two Models in Terms of Taylor Approximation
The loss function $\mathcal{L}:U\ni\bm{\theta}\mapsto\mathcal{L}(\bm{\theta})\in\mathbb{R}$ is assumed to be a twice differentiable function on an open set $U$ over $\mathbb{R}^{d_{\textrm{param}}}$. Let $\bm{H}_{a}$ and $\bm{H}_{b}$ be the Hessian matrices centered at the models $\bm{\theta}_{a}$ and $\pi(\bm{\theta}_{b})$, respectively. If, for any $\lambda\in(0,1)$, $\lambda\bm{\theta}_{a}+(1-\lambda)\bm{\theta}_{b}\in U$ holds, then we have
@@
where $\nabla$ is the gradient with respect to the parameters, $O$ is the Landau symbol, $\beta$ is the $L_{2}$ distance between $\bm{\theta}_{a}$ and $\pi(\bm{\theta}_{b})$, and $\bm{\mu}$ is the unit vector from $\bm{\theta}_{a}$ to $\pi(\bm{\theta}_{b})$ (i.e., $\bm{\mu}=(\pi(\bm{\theta}_{b})-\bm{\theta}_{a})/\beta$).
@@ 3.2 Experimental Results
Table 1: Results of WM and the estimated barrier value using Taylor approximation when $\lambda=1/2$. The table presents the mean and standard deviation from five trials of model merging (i.e., the linear combination of the models $(\bm{\theta}_{a}+\pi(\bm{\theta}_{b}))/2$). The columns labeled “Test acc.”, “Barrier”, and “Taylor approx.” show the test accuracy, the barrier value, and the estimated barrier value using Equation 3 for the merged model at $\lambda=1/2$, respectively. The table also shows the $L_{2}$ distance between the models $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$ before and after applying the permutation, as well as the reduction rate of the $L_{2}$ distance (i.e., $(\|\bm{\theta}_{a}-\bm{\theta}_{b}\|-\|\bm{\theta}_{a}-\pi(\bm{\theta}_{b})\|)/\|\bm{\theta}_{a}-\bm{\theta}_{b}\|$).
CIFAR10 VGG11 $85.456\pm 3.260$ $0.035\pm 0.1$ $2.956\pm 0.35$ $799.503\pm 16.396$ $746.465\pm 19.576$ $6.64\pm 0.808$
ResNet20 $88.808\pm 0.369$ $0.167\pm 0.035$ $7.517\pm 0.573$ $710.762\pm 16.261$ $661.055\pm 12.539$ $6.987\pm 0.472$
FMNIST MLP $88.330\pm 0.438$ $-0.183\pm 0.049$ $0.928\pm 0.175$ $121.853\pm 5.83$ $100.041\pm 4.71$ $17.897\pm 0.348$
MNIST MLP $98.382\pm 0.116$ $-0.033\pm 0.006$ $0.036\pm 0.03$ $81.231\pm 5.58$ $64.751\pm 4.795$ $20.305\pm 1.225$
We conducted experiments to verify whether the Taylor approximation (Equation 3) accurately estimates the barrier between two SGD solutions. Table 1 presents the experimental results of model merging. Details about the datasets, network training procedures, and permutation search methods used in these experiments are described in Appendix D. In the table, we chose $\lambda=1/2$ because it is empirically known that the midpoint between two models results in the highest loss (Ainsworth et al., 2023; Peña et al., 2023). We used an algorithm similar to the backpropagation method (Bishop, 2006; Pearlmutter, 1994) to efficiently compute $\bm{\mu}^{\top}\bm{H}$, which is required for the evaluation of Equation 3. A negative value in the Barrier column indicates that the loss of the merged model is lower than that of the original (pre-merged) models.
The table shows that, except for the MNIST dataset, there is a large difference between the actual barrier values and those estimated by the Taylor approximation. These differences are particularly large for VGG11 and ResNet20. Additionally, the table indicates that the $L_{2}$ distance changes by only about 6% to 20% from the original distance. This suggests that WM does not bring the models sufficiently close, at least not close enough for a second-order Taylor approximation to hold.

### v5

@@ Abstract
Recently, Ainsworth et al. (2023) showed that using weight matching (WM) to minimize the $L^{2}$ distance in a permutation search of model parameters effectively identifies permutations that satisfy linear mode connectivity (LMC), where the loss along a linear path between two independently trained models with different seeds remains nearly constant. This paper analyzes LMC using WM, which is useful for understanding stochastic gradient descent’s effectiveness and its application in areas like model merging. We first empirically show that permutations found by WM do not significantly reduce the $L^{2}$ distance between two models, and the occurrence of LMC is not merely due to distance reduction by WM itself. We then demonstrate that permutations can change the directions of the singular vectors, but not the singular values, of the weight matrices in each layer. This finding shows that permutations found by WM primarily align the directions of singular vectors associated with large singular values across models. This alignment brings the singular vectors with large singular values, which determine the model’s functionality, closer between the original and merged models, allowing the merged model to retain functionality similar to the original models, thereby satisfying LMC. This paper also analyzes activation matching (AM) in terms of singular vectors and finds that the principle of AM is likely the same as that of WM. Finally, we analyze the difference between WM and the straight-through estimator (STE), a dataset-dependent permutation search method, and show that WM can be more advantageous than STE in achieving LMC among three or more models.
@@ 1 Introduction
Large-scale neural networks (NNs) are widely used in various fields (Vaswani et al., 2017; van den Oord et al., 2016; Zhao et al., 2023), and optimizing their parameters poses a massive non-convex optimization problem. Remarkably, stochastic gradient descent (SGD), which is widely used for training NNs, is known to find good solutions despite its simplicity. One hypothesis for this seemingly counterintuitive phenomenon is that the landscape of the loss function may be much simpler than previously thought. Several studies (Garipov et al., 2018; Draxler et al., 2018; Freeman & Bruna, 2017) have found that different NN solutions can be connected by simple nonlinear paths with almost no increase in loss. Recently, Entezari et al. (2022) conjectured that 1.1 holds, considering all possible permutation symmetries of NNs:
@@
Here, the barrier represents the increase in loss observed when linearly interpolating between the weights of the two models. If the barrier between two models is sufficiently small, we say that linear mode connectivity (LMC) is satisfied between them (Frankle et al., 2020). 1.1 suggests that most SGD solutions can be transferred into the same loss basin using permutations. Indeed, some studies (Ainsworth et al., 2023; Singh & Jaggi, 2020) have experimentally demonstrated that this conjecture is valid for various datasets and models using weight matching (WM), which identifies permutations that minimize the $L^{2}$ distance between the weights of two models.
Understanding LMC principles based on permutation symmetries is important not only for comprehending how SGD works in deep learning but also for its application in model merging (Singh & Jaggi, 2020), where two independently trained models are combined. The method of finding permutations using only the $L^{2}$ distance is particularly versatile, dataset-independent, and computationally efficient. In fact, several studies (Singh & Jaggi, 2020; Wang et al., 2020; Guerrero-Peña et al., 2023) have proposed applications of permutation symmetries in model merging, federated learning, and continual learning.
The current theoretical analysis of LMC relies on the feasibility of closely matching NN weights through permutations. Recently, Zhou et al. (2023) proved that if the distance between the weights of two models can be sufficiently reduced via permutation, then LMC holds. Intuitively, for two SGD solutions $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$, if $\bm{\theta}_{a}\approx\pi(\bm{\theta}_{b})$ holds for a permutation $\pi$, the outputs of the interpolated model will closely approximate those of the original models $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$.
However, our analysis reveals that even if LMC holds, the permutations found by WM do not significantly reduce the distance between the two models (at most about a 20% reduction). This suggests that LMC is satisfied even when WM does not bring the two models very close (i.e., $\bm{\theta}_{a}\not\approx\pi(\bm{\theta}_{b})$). Accordingly, this paper seeks to uncover a more fundamental reason why LMC holds through the permutations found by WM. Specifically, we demonstrate that singular vectors with large singular values of each weight in the models play a crucial role in LMC. Our analysis not only reveals the principle behind WM but also shows that WM may be more advantageous in merging more than two models compared to other methods such as STE.
@@
1. Demonstrating that the $L^{2}$ distance reduced by WM is not the direct cause of LMC.
We empirically show that permutations found by WM do not significantly reduce the $L^{2}$ distance between the two models. Our results that, even when LMC is satisfied, permutations reduce the model weight distance by no more than 20%. Supported by a Taylor approximation, our findings suggest that reducing the $L^{2}$ distance through permutations is not the direct reason for LMC satisfaction.
@@
We analyze WM from the perspective of the function of each layer of the model. Specifically, we provide evidence that WM satisfies LMC by aligning the directions of singular vectors with large singular values in each layer’s weights. This alignment ensures that the singular vectors with large singular values, which determine the functionality of each layer, become similar between the merged and original models. Additionally, we show that, from the perspective of the input distribution at each hidden layer, aligning singular vectors with large singular values can efficiently approximate the functionality between two models, even if the $L^{2}$ distance cannot be significantly reduced. As a result, the merged model retains functionality similar to the original models, which facilitates LMC. We also conducted experiments with AM and found that the reason why LMC holds in AM is likely the same as in WM.
@@
To distinguish WM from other permutation search methods that are independent of $L^{2}$ distance, we examine the straight-through estimator (STE), which focuses on minimizing the barrier itself rather than the $L^{2}$ distance. Our experiments reveal that the permutations found by STE do not align the directions of singular vectors, which is a critical difference compared to WM in achieving LMC. Furthermore, we demonstrate experimentally that this difference significantly impacts the satisfaction of LMC among three or more models.
@@ 2.1 Notation
For any natural number $k\in\mathbb{N}$, let $[k]=\{1,2,\dots,k\}$. Bold uppercase variables represent tensors, including matrices (e.g., $\bm{X}$), and bold lowercase variables (e.g., $\bm{x}$) represent vectors. For any tensor $\bm{X}$, its vectorization is denoted by $\textrm{vec}(\bm{X})$, and $\|\bm{X}\|$ denotes its Frobenius ($L^{2}$) norm.
@@ 2.3 Linear Mode Connectivity (LMC)
Intuitively, the barrier represents the increase in loss due to the linear interpolation of the two models. Two models $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$ are said to be linearly mode connected if their loss barrier is approximately zero.
@@ 2.4 Permutation Selection
In WM, we search for a permutation that minimizes the $L^{2}$ distance between two models [^1] Although only the weights are considered here, the biases can also be dealt with by concatenating the biases and the weights.:
@@
where, without loss of generality, let $\bm{P}_{L}=\bm{I}$ and $\bm{P}_{0}=\bm{I}$, and $\bm{I}$ is an identity matrix. This minimization problem is known as the sum of the bilinear assignments problem, which is NP-hard (Koopmans & Beckmann, 1957; Sahni & Gonzalez, 1976; Ainsworth et al., 2023). Recently, Guerrero-Peña et al. (2023) proposed solving Equation 2 using Sinkhorn’s algorithm (Adams & Zemel, 2011) by using an optimal transport problem. We adopt their method because it allows for the optimization of all layers simultaneously, unlike the method by Ainsworth et al. (2023), and potentially finds better solutions.
@@ 3 Motivating Observations
Previous studies have suggested that the closeness of two parameters in terms of $L^{2}$ distance is important for satisfying LMC. For example, Zhou et al. (2023) showed that LMC holds if a commutativity property is satisfied. This property holds if, for all layers $\ell$, $\bm{W}^{(a)}_{\ell}-\bm{P}_{\ell}\bm{W}^{(b)}_{\ell}\bm{P}_{\ell-1}^{\top}=\bm{0}$. Zhou et al. (2023) argued in Section 5.2 that since WM finds the permutation that minimizes Equation 2, WM can be seen as searching for permutations that satisfy the commutativity property. In particular, there is a huge number of permutations because the total number of possible permutations grows exponentially as the number of layers and the width increase, and thus, some of them may sufficiently reduce the distance between the two models. However, this section explains from the perspective of a Taylor approximation that this intuition is not always correct. Our results demonstrate that even when LMC is satisfied, the permutations found by WM do not necessarily bring the models as close as expected. The facts observed from the experiments in this section motivate us to explore other reasons for satisfying LMC in the following sections.
@@ 3.1 Closeness of Two Models in Terms of Taylor Approximation
The loss function $\mathcal{L}:U\ni\bm{\theta}\mapsto\mathcal{L}(\bm{\theta})\in\mathbb{R}$ is assumed to be of class $C^{3}$ on an open set $U$ over $\mathbb{R}^{d_{\textrm{param}}}$. Let $\bm{H}_{a}$ and $\bm{H}_{b}$ be the Hessian matrices centered at the models $\bm{\theta}_{a}$ and $\pi(\bm{\theta}_{b})$, respectively. If, for any $\lambda\in(0,1)$, $\lambda\bm{\theta}_{a}+(1-\lambda)\bm{\theta}_{b}\in U$ holds, then we have
@@
where $\nabla$ is the gradient with respect to the parameters, $O$ is the Landau symbol, $\beta$ is the $L^{2}$ distance between $\bm{\theta}_{a}$ and $\pi(\bm{\theta}_{b})$, and $\bm{\mu}$ is the unit vector from $\bm{\theta}_{a}$ to $\pi(\bm{\theta}_{b})$ (i.e., $\bm{\mu}=(\pi(\bm{\theta}_{b})-\bm{\theta}_{a})/\beta$).
@@ 3.2 Experimental Results
Table 1: Results of WM and the estimated barrier value using Taylor approximation when $\lambda=1/2$. The table presents the mean and standard deviation from five trials of model merging (i.e., the linear combination of the models $(\bm{\theta}_{a}+\pi(\bm{\theta}_{b}))/2$). The columns labeled “Barrier”, “Taylor approx.”, and “Diff.” show the barrier value, the estimated barrier value using Equation 3 for the merged model at $\lambda=1/2$, and their difference, respectively. In the “Diff.” column, if a statistical significant difference is determined using a t-test at a 5% significance level, they are highlighted in bold. The table also shows the $L^{2}$ distance between the models $\bm{\theta}_{a}$ and $\bm{\theta}_{b}$ before and after applying the permutation, as well as the reduction rate of the $L^{2}$ distance (i.e., $(\|\bm{\theta}_{a}-\bm{\theta}_{b}\|-\|\bm{\theta}_{a}-\pi(\bm{\theta}_{b})\|)/\|\bm{\theta}_{a}-\bm{\theta}_{b}\|$).
Dataset Network Barrier ($\lambda=1/2$) Taylor approx. Diff. $\|\bm{\theta}_{a}-\bm{\theta}_{b}\|$ $\|\bm{\theta}_{a}-\pi(\bm{\theta}_{b})\|$ Reduction rate [%]
CIFAR10 VGG11 $0.035\pm 0.1$ $2.956\pm 0.35$ $\mathbf{2.921\pm 0.323}$ $799.503\pm 16.396$ $746.465\pm 19.576$ $6.64\pm 0.808$
ResNet20 $0.167\pm 0.035$ $7.517\pm 0.573$ $\mathbf{7.349\pm 0.599}$ $710.762\pm 16.261$ $661.055\pm 12.539$ $6.987\pm 0.472$
FMNIST MLP $-0.183\pm 0.049$ $0.928\pm 0.175$ $\mathbf{1.111\pm 0.152}$ $121.853\pm 5.83$ $100.041\pm 4.71$ $17.897\pm 0.348$
MNIST MLP $-0.033\pm 0.006$ $0.036\pm 0.03$ $\mathbf{0.069\pm 0.028}$ $81.231\pm 5.58$ $64.751\pm 4.795$ $20.305\pm 1.225$
We conducted experiments to verify whether the Taylor approximation (Equation 3) accurately estimates the barrier between two SGD solutions. Table 1 presents the experimental results of model merging. Details about the datasets, network training procedures, and permutation search methods used in these experiments are described in Appendix D. In the table, we chose $\lambda=1/2$ because it is empirically known that the midpoint between two models results in the highest loss (Ainsworth et al., 2023; Guerrero-Peña et al., 2023). It is worth noting that Adilova et al. (2024) demonstrated that the location of the highest barrier shifts when one model is more generalized than the other, although not directly applicable to this paper as the two models are generalized to the same extent. We used the vhp function provided in the PyTroch library [^2] https://pytorch.org/docs/stable/generated/torch.autograd.functional.vhp.html to efficiently compute $\bm{\mu}^{\top}\bm{H}$, which is required for the evaluation of Equation 3. A negative value in the Barrier column indicates that the loss of the merged model is lower than those of the original (pre-merged) models.
The table shows that for all datasets, there is a significant difference between the actual barrier values and those estimated by the Taylor approximation. These differences are particularly large for VGG11 and ResNet20. Additionally, the table indicates that the $L^{2}$ distance changes by only about 6% to 20% from the original distance. This suggests that WM does not bring the models sufficiently close, at least not close enough for a second-order Taylor approximation to hold.